# MimicTouch: Leveraging Multi-modal Human Tactile Demonstrations for Contact-rich Manipulation

### v3

@@ Abstract
Tactile sensing is critical to fine-grained, contact-rich manipulation tasks, such as insertion and assembly. Prior research has shown the possibility of learning tactile-guided policy from teleoperated demonstration data. However, to provide the demonstration, human users often rely on visual feedback to control the robot. This creates a gap between the sensing modality used for controlling the robot (visual) and the modality of interest (tactile). To bridge this gap, we introduce “MimicTouch”, a novel framework for learning policies directly from demonstrations provided by human users with their hands. The key innovations are i) a human tactile data collection system which collects multi-modal tactile dataset for learning human’s tactile-guided control strategy, ii) an imitation learning-based framework for learning human’s tactile-guided control strategy through such data, and iii) an online residual RL framework to bridge the embodiment gap between the human hand and the robot gripper. Through comprehensive experiments, we highlight the efficacy of utilizing human’s tactile-guided control strategy to resolve contact-rich manipulation tasks. The project website is at https://sites.google.com/view/MimicTouch.
Figure 1: The first row shows the human tactile demonstrations, including the tactile and proprioception data. The second row shows the robot execution with tactile feedback. The third row below the dashed line describes the policy’s zero-shot generalization capability in five different domains, including variations in hole positions, angles, inner shapes, materials, and a different assembly task
@@ 2 Related Works
Multi-modal tactile sensing. Vision-based tactile sensing is integral to robotic manipulation as it excels at estimating local contact geometry and frictional properties [13, 14, 15, 16]. It further enables imitation learning-based methods through policy observations [6, 17, 7] or reinforcement learning methods through reward signals [18, 19, 20] for various complex manipulation tasks. Additionally, it is also widely used in shape reconstruction [21, 22, 23] and grasping [24, 25, 26, 27]. On the other hand, audio-based tactile sensors, such as contact microphones have also been demonstrated effective in robotics applications such as manipulation [28], classification [29, 30], and dynamics modelling [31, 32]. These sensors can emulate nerve endings within human skin to better detect vibrations during tactile interactions. Therefore, incorporating both sensor modalities can yield tactile feedback more akin to human sensations, enabling the robot to learn a human-like tactile-guided control strategy.
@@ 3 MimicTouch Framework
We aim to enable the robot to resolve contact-rich manipulation tasks by learning control strategy from human tactile demonstrations. To achieve this, we propose a novel learning framework named “MimicTouch”. It first introduces a human tactile-guided data collection system (Sec. 3.1) to collect a multi-modal tactile dataset from human demonstrators. Then, to emulate the human’s tactile-guided control strategy for successful robot execution, MimicTouch has three distinct learning phases. Firstly, it learns lower dimensional tactile representations from the human tactile demonstrations in a self-supervised manner (Sec. 3.2). Next, it derives an offline policy with the learned representations using a non-parametric imitation learning method [12] (Sec. 3.3). Lastly, it refines the offline policy through online residual reinforcement learning (Sec. 3.4). Note that this refinement phase is efficient and reliable as the offline policy encodes human’s tactile-guided control strategy. The overall MimicTouch framework is shown in Fig. 2.
@@ 3.2 Learning Tactile Representation
The policy learning on high-dimensional sensor inputs struggles with real-world deployment due to computational burden and sensor noise. Additionally, in this work, variations may exist between sensor inputs from human tactile demonstrations and real-time robot sensor feedback due to discrepancies in finger-object contact force. Therefore, inspired by recent works that learn lower-dimensional embeddings for image-guided imitation learning [6, 48, 12] which can discover the appropriate features that are helpful for policy learning, we first learn the compact representation for both tactile and audio sensor data using self-supervised learning methods (part (b) in Fig. 2). Intuitively, it identifies a low-dimensional embedding space where differently augmented images, such as tactile images or audio spectrum, are projected to a similar embedding. As a result, these embeddings are more efficient for online processing and more robust to task-irrelevant sensor noise. This learning phase consists of the following two parts.
@@
Self-supervised learning. We employ the Bootstrap Your Own Latent (BYOL) [49] for tactile images and BYOL for audio (BYOL-A) for audio segments [50], since they have demonstrated desired performance in computer vision [49], audio representation [50], and robotics [7, 36, 28] tasks. For BYOL and BYOL-A, the inputs are tactile images and audio spectrums separately, and the embedding spaces are both configured to be 2048-dimensional. Learning details are included in the Appendix. B.
@@ 3.3 Learning Offline Policy from Human Tactile Demonstrations
Leveraging the learned representations, we then learn the robot policy from the human tactile demonstrations. Here, one unique challenge is that the human hand moves much faster than the robot, resulting in sparse temporal observation-action samples (i.e., large action values per observation). Also, human tactile demonstrations might partially be out-of-domain (OOD) demonstrations for robot policy due to the embodiment gap (e.g., different motion capability and finger-object contact forces). Therefore, executing a parametric policy might exhibit unreasonable robot behaviors as they are prone to covariant shift [44] (see Sec. 4.2.1 for experimental validation). As a result, we use a non-parametric imitation learning method [12] to ensure the execution efficacy of the policy learned from human tactile demonstrations (part (c) in Fig. 2). In addition to the learning algorithm, data pre-processing is necessary for synchronization and the details are included in Appendix. C.
Non-parametric imitation learning. We build our algorithm on the VINN framework [12], extending it to include tactile-audio representation without relying on visual signals. At the $i$-th time step, the observations and actions are denoted as $(\mathrm{o}^{T}_{i},\mathrm{o}^{A}_{i},\mathrm{o}^{EE},\mathrm{a}_{i})$, where $\mathrm{o}^{T}$ is the tactile representation, $\mathrm{o}^{A}$ is the audio representation, $\mathrm{o}^{EE}$ is the robot end-effector pose, and the action $\mathrm{a}$ is defined by the $6D$ delta pose of the robot end-effector, including a delta position and a delta Euler angle. Then, we extract tactile and audio features $(\mathrm{y}^{T}_{i},\mathrm{y}^{A}_{i})$ from $(\mathrm{o}^{T}_{i},\mathrm{o}^{A}_{i})$ using the pre-trained representation encoders, respectively. These tactile embeddings and the robot end-effector pose $(\mathrm{y}^{T}_{i},\mathrm{y}^{A}_{i},\mathrm{o}^{EE}_{i})$ are formulated as the key features of the demonstration library, with each associated with a corresponding action value $\mathrm{a}_{i}$. Given the varying scales of these inputs, we normalize them such that the maximum distance for each input is unity in the library. In robot execution, for a given real-time observation  $(\hat{\mathrm{o}}^{T}_{i},\hat{\mathrm{o}}^{A}_{i},\hat{\mathrm{o}}^{EE}_{i})$, we first obtain the query feature $(\hat{\mathrm{y}}^{T}_{t},\hat{\mathrm{y}}^{A}_{t},\hat{\mathrm{o}}^{EE}_{t})$, and then search the demonstration library for a nearest-neighbor-based action prediction.
@@ 3.4 Learning Residual Policy through Online Reinforcement Learning
The offline policy learned from human tactile demonstrations might not guarantee task success when deployed on the physical robot. This could be due to: i). morphological differences between the human hand and the robot gripper, ii). inaccurate fingertip tracking caused by fast movements, and iii). underexplored contact effects. Therefore, motivated by recent works using pure RL [17, 18, 19, 20, 51] to learn tactile policies, we further leverage online reinforcement learning that allows in-domain robot interactions (part (d) of Fig. 2). It is noteworthy that the previous pure RL methods often generate quasi-static motions and utilize a limited action space [17, 18, 19, 51] because they learn from scratch without effective priors. On the contrary, we intend to leverage the best of both advantages by RL fine-tuning the offline policy learned from human tactile demonstrations.
Since it is infeasible to directly fine-tune the non-parametric policy, we instead learn a residual policy. The input to the residual policy $\pi_{r}$ consists of the current observation $(\hat{\mathrm{y}}^{T}_{i},\hat{\mathrm{y}}^{A}_{i},\hat{\mathrm{o}}^{EE}_{i})$ and the offline policy’s output $a_{i}^{o}$ and the output is the residual action $a_{i}^{r}$. Considering we use $6D$ continuous action space and sparse observation-action pairs (around 70 actions per trajectory), we opt for SAC [52] to handle the continuous action space with entropy regularization and to generate a replay buffer to increase the size of training data. Finally, the robot action is the sum of the action generated from offline policy $\pi_{o}$ and the action from the residual policy $\pi_{r}$. To ensure that the residual policy only makes slight adjustments, we set specific action limits for the residual policy.
Another critical component of residual RL policy learning is the reward design, which must balance exploitation and exploration. To address this, we combine an expert-aligned reward, which is measured by the KL divergence between the human expert trajectories and the robot executed trajectory, with a task-specific reward. The expert-aligned reward encourages a policy distribution that mimics the demonstrations, whereas the task-specific reward drives exploration to optimize the in-domain robot policy. More details about pseudocode, action limits, policy design, and rewards are included in Appendix. D
@@ 4 Experiments
Teleoperation setting. We compare our human tactile-guided data collection system (Sec. 3.1) with Spacemouse-based teleoperation, a commonly used interface for manipulation tasks [9, 11, 37], and a Hand-guided teleoperation (See Appendix. E for details. For both, to collect a similar number of observation-action pairs for each trajectory, we collect one robot state, one tactile image, and 0.5s audio segment at 5Hz. Since Spacemouse-based teleoperation requires considerable expertise, we allocate approximately 5 hours for practice with this system. In contrast, Hand-guided teleoperation, similar to our proposed system, does not require any practice.
Tasks. We focus on two-piece insertion tasks that exemplify the challenge of many contact-rich manipulation tasks. We 3D-print a cylinder and an insertion hole base and set up the same task environments for both data collection settings. An example has been shown in Fig. 1, and the dimensions of each piece, including those used in the generalization tasks (Sec. 4.3), are detailed in Appendix. F.
@@ 4.1 Human Tactile Demonstration Collection System
The results in Table. 1 support our insights: i) human tactile demonstrations can be collected significantly more efficiently than teleoperation systems for contact-rich tasks, and ii) human tactile demonstrations can seamlessly integrate human’s tactile feedback and motion capability, whereas teleoperation systems struggle to capture such dynamic tactile-guided motions. These factors together result in much lower data collection efficiency and success rates of the teleoperation systems.
@@ 4.2.1 MimicTouch effectively learns from human tactile demonstrations
In this subsection, we evaluate the offline policy learned from the human tactile demonstrations using both the nearest-neighbor-based method and a parametric imitation learning method. Specifically, we aim to test whether the offline policies can be deployed in real-world environments within the desired error tolerance, which is crucial for physical robot execution. To evaluate it, we use the data collection system (Sec. 3.1) to collect 20 noise-free data sequences as the datasets to learn both offline policies. For testing, we first gather another 5 data sequences with random noise to emulate the real-world environments (details are in Appendix. G), and compute the mean square error loss (MSE Loss, defined in Appendix. H) on the testing set. Also, we conduct the real-world experiments and compare the task success rate over 25 policy rollouts.
@@
Additionally, we conduct the ablation study on the choice of multi-modal sensor inputs in Appendix. J. The results suggest that multi-modal tactile feedback is crucial for the success of contact-rich insertion tasks, particularly during the insertion phase when most contact occurs.
@@ 4.2.2 Human tactile demonstrations trains better policies than teleoperated demonstrations
We first report that the task success rate of the offline policy rollouts from human tactile demonstration is 40% (10/25), whereas the success rate from Spacemouse-based teleoperation rollouts is only 12% (3/25), and from Hand-guided teleoperation, it is 28% (7/25). Then, in the bottom part of Fig. 3, we show the mean and variance of the action serial numbers of three successful rollout trajectories for each offline policy. For the human tactile demonstrations, we observe a linear relationship with minimal variance. On the contrary, the action serial numbers for the both teleoperation policies exhibit greater nonlinearity with significantly larger variance, particularly during the insertion phase. This performance discrepancy arises because the majority of contacts occur during this phase, and the teleoperation systems that lack of human tactile feedback is not well-suited for capturing these contact-rich events. Therefore, the indexes of the selected key features are more disordered.
@@
A similar conclusion can be drawn from the screenshots (top part of Fig. 3) for one successful rollout of all policies. The Human Tactile Demonstrations policy exhibits dynamic tilting for object insertion, which emulates the demonstrated human’s tactile-guided control strategy (see Fig. 1). However, the teleoperation policies forcefully inserts the object with less orientation, indicating that teleoperation demonstrations capture less versatile motions that are necessary for contact-rich tasks.
@@ 4.3 Online Policy Improvement and Generalization to New Settings
In this section, we evaluate the final policy trained through online reinforcement learning (RL). To ensure the robustness of the RL residual policy, we perform domain randomization on the robot’s starting position so that the initial object-hole contact is located differently. For the policy update, at each iteration, we use five newly collected trajectories along with another five randomly selected trajectories from the replay buffer.
@@
MimicTouch policy exhibits superior zero-shot generalization capability. We evaluate the zero-shot generalizability of the MimicTouch final policy. We consider the following generalization settings: i). Shifting Positions: an insertion task with the hole positions shifted for 0.8cm in either ±$\pm x$ or ±$\pm y$ horizontal directions, ii). Tilting Angles: an insertion task with the hole angle tilted for 10 degrees or 20 degrees, iii). Two-stage Dense Packing: a two-stage dense insertion task, which requires the robot to perform two consecutive insertion alignments, iv). Multi-material Dense Packing: an insertion task where the hole contains multiple objects, such as a pen, tissues, or a sponge, and v) Furniture Assembly [53, 54]: an insertion task, which requires the robot to insert and adjust the leg into a small hole in a table for screwing. Each setting is depicted in Fig. 5. Additionally, to demonstrate the complexity of these tasks, we introduce another baseline: Openloop Policy, where we collect five successful insertion trajectories in the initial setting and execute each of these trajectories for those generalization tasks five times for each of the tasks. See Appendix. L for the details of each task and the policy evaluation process.
@@
Table 2: Task success rate for each generalization task. Specifically, Shift is Shifting Position, 10°and 20°are different tilted angles for Tilting Angles, Two-stage is two-stage Dense Packing, Rigid means the object contacts the pen in Multi-material Dense Packing, Soft means the object contacts the tissue and sponge in Multi-material Dense Packing, Assem (I), and Assem are the insertion-only, and insertion-adjustment results for Furniture Assembly respectively.

### v4

@@ Abstract
Tactile sensing is critical to fine-grained, contact-rich manipulation tasks, such as insertion and assembly. Prior research has shown the possibility of learning tactile-guided policy from teleoperated demonstration data. However, to provide the demonstration, human demonstrators often rely on visual feedback to control the robot. This creates a gap between the sensing modality used for controlling the robot (visual) and the modality of interest (tactile). To bridge this gap, we introduce “MimicTouch”, a novel framework for learning policies directly from demonstrations provided by human users with their hands. The key innovations are i) a human tactile data collection system which collects multi-modal tactile dataset for learning human’s tactile-guided control strategy, ii) an imitation learning-based framework for learning human’s tactile-guided control strategy through such data, and iii) an online residual RL framework to bridge the embodiment gap between the human hand and the robot gripper. Through comprehensive experiments, we highlight the efficacy of utilizing human’s tactile-guided control strategy to resolve contact-rich manipulation tasks. The project website is at https://sites.google.com/view/MimicTouch.
Figure 1: The first row shows the human tactile demonstrations, including the tactile and proprioception data. The second row shows the robot execution with tactile feedback. The third row below the dashed line describes the policy’s zero-shot generalization capability in five different domains, including variations in hole positions, angles, inner shapes, materials, and a different assembly task.
@@ 2 Related Works
Multi-modal tactile sensing. Vision-based tactile sensing is integral to robotic manipulation as it excels at estimating local contact geometry and frictional properties [13, 14, 15, 16]. It further enables imitation learning-based methods through policy observations [6, 17, 7] or reinforcement learning methods through reward signals [18, 19, 20] for various contact-rich manipulation tasks. Additionally, it is also widely used in shape reconstruction [21, 22, 23] and grasping [24, 25, 26, 27]. On the other hand, audio-based tactile sensors, such as contact microphones have also been demonstrated effective in robotics applications such as manipulation [28], classification [29, 30], and dynamics modelling [31, 32]. These sensors can emulate nerve endings within human skin to better detect vibrations during tactile interactions. Therefore, incorporating both sensor modalities can yield tactile feedback more akin to human sensations, enabling the robot to learn a human-like tactile-guided control strategy.
@@ 3 MimicTouch Framework
We aim to enable the robot to resolve contact-rich manipulation tasks by learning tactile-guided control strategy from human demonstrations. To achieve this, we propose a novel learning framework named “MimicTouch”. It first introduces human tactile demonstrations (Sec. 3.1) collected by human hands. Then, to emulate the human’s tactile-guided control strategy for successful robot execution, MimicTouch has three learning phases. Firstly, it learns lower dimensional tactile representations from the human tactile demonstrations in a self-supervised manner (Sec. 3.2). Next, it derives an offline policy with the learned representations using a non-parametric imitation learning method [12] (Sec. 3.3). Lastly, it refines the offline policy through online residual RL on the real robot(Sec. 3.4). The overall MimicTouch framework is shown in Fig. 2.
@@ 3.2 Learning Tactile Representation
The policy learning on high-dimensional sensor inputs struggles with real-world deployment due to computational burden and sensor noise. Additionally, in this work, variations may arise between sensor inputs from human tactile demonstrations and real-time robot feedback due to different finger-object contact force. Inspired by recent works that learn lower-dimensional embeddings for imitation learning [6, 48, 12], we learn the compact representation for both tactile and audio data using self-supervised learning methods (part (b) in Fig. 2). Intuitively, it identifies a low-dimensional embedding space for different augmented data (tactile images or audio spectrum) and projects them to a similar embedding. As a result, these embeddings are more computationally efficient and more robust to task-irrelevant sensor noise. This learning phase consists of the following two parts:
@@
Self-supervised learning. We employ the Bootstrap Your Own Latent (BYOL) [49] for tactile images and BYOL for audio (BYOL-A) for audio segments [50], since they have demonstrated desired performance in computer vision [49], audio learning [50], and robotics [7, 36, 28]. For BYOL and BYOL-A, the inputs are tactile images and audio spectrums separately, and the embedding spaces are both configured to be 2048-dimensional. Learning details are included in the Appendix. B.
@@ 3.3 Learning Offline Policy from Human Tactile Demonstrations
The next step is learning robot policy from the human tactile demonstrations. Here, one unique challenge is that the human hand moves much faster than the robot, resulting in sparse temporal observation-action samples (i.e., large action values per observation). Also, the embodiment gap (e.g., different motion capabilities and contact forces) can lead to out-of-domain demonstrations, which will make parametric policies prone to covariant shift [44] (see Sec. 4.2.1 for experimental validation). As a result, we use a non-parametric imitation learning method [12] to ensure the execution efficacy of the policy learned from human tactile demonstrations (part (c) in Fig. 2). In addition to the learning algorithm, data pre-processing is necessary for synchronization and the details are included in Appendix. C.
Non-parametric imitation learning. We build our algorithm on the NN-based framework [12], extending it to include tactile-audio representation without visual signals. At the $i$-th time step, the observations and actions are denoted as $(\mathrm{o}^{T}_{i},\mathrm{o}^{A}_{i},\mathrm{o}^{EE},\mathrm{a}_{i})$, where $\mathrm{o}^{T}$ is the tactile representation, $\mathrm{o}^{A}$ is the audio representation, $\mathrm{o}^{EE}$ is the robot end-effector pose, and the action $\mathrm{a}$ is defined by the $6D$ delta pose of the robot end-effector. Then, we extract tactile and audio features $(\mathrm{y}^{T}_{i},\mathrm{y}^{A}_{i})$ from $(\mathrm{o}^{T}_{i},\mathrm{o}^{A}_{i})$ using the pre-trained representation encoders, respectively. These tactile embeddings and the robot end-effector pose $(\mathrm{y}^{T}_{i},\mathrm{y}^{A}_{i},\mathrm{o}^{EE}_{i})$ are formulated as the key features of the demonstration library. Given the varying scales of these inputs, we normalize them such that the maximum distance for each input is unity in the library. In robot execution, for a given real-time observation  $(\hat{\mathrm{o}}^{T}_{i},\hat{\mathrm{o}}^{A}_{i},\hat{\mathrm{o}}^{EE}_{i})$, we first obtain the query feature $(\hat{\mathrm{y}}^{T}_{t},\hat{\mathrm{y}}^{A}_{t},\hat{\mathrm{o}}^{EE}_{t})$, and then search the demonstration library for a nearest-neighbor-based action prediction.
@@ 3.4 Learning Residual Policy through Online Reinforcement Learning
The offline policy learned from human tactile demonstrations might not guarantee task success when deployed on the physical robot. This could be due to: i). morphological differences between the human hand and the robot gripper, ii). inaccurate fingertip tracking caused by fast movements, and iii). underexplored contact effects. Therefore, motivated by recent works using pure RL [17, 18, 19, 20, 51] to learn tactile policies, we further leverage online reinforcement learning that allows in-domain robot interactions (part (d) of Fig. 2). It is noteworthy that the previous pure RL methods often generate quasi-static motions and utilize a limited action space [17, 18, 19, 51] because they learn from scratch without effective motion priors. On the contrary, we intend to leverage the best of both advantages by RL fine-tuning the offline policy learned from human tactile demonstrations.
Since it is infeasible to directly fine-tune the non-parametric policy, we instead learn a residual policy. The input to the residual policy $\pi_{r}$ consists of the current observation $(\hat{\mathrm{y}}^{T}_{i},\hat{\mathrm{y}}^{A}_{i},\hat{\mathrm{o}}^{EE}_{i})$ the last policy output, which is the sum of the offline policy output and the residual policy output $a_{i-1}^{o}+a_{i-1}^{r}$. Considering we use $6D$ continuous action space and sparse observation-action pairs (around 70 actions per trajectory), we opt for SAC [52] to handle the continuous action space with entropy regularization and to generate a replay buffer to increase the size of training data. Finally, the robot action is the su, of the offline policy output and the residual policy output. To ensure that the residual policy only makes slight adjustments, we set specific action limits for the residual policy.
Another critical component of residual RL is the reward design, which must balance exploitation and exploration. To address this, we combine an expert-aligned reward, which is measured by the KL divergence between the human expert trajectory and the robot executed trajectory, with a task-specific reward. The expert-aligned reward encourages a policy distribution that mimics the demonstrations, whereas the task-specific reward drives exploration to optimize the in-domain robot policy. More details about pseudocode, action and policy design, and rewards are included in Appendix. D.
@@ 4 Experiments
Teleoperation setting. We compare our human tactile-guided data collection system (Sec. 3.1) with Spacemouse-based teleoperation, a commonly used interface for manipulation tasks [9, 11, 37], and a Hand-guided teleoperation similar to Kinesthetic teaching (See Appendix. E for details). For both, to collect a similar number of observation-action pairs for each trajectory, we collect one robot state, one tactile image, and 0.5s audio segment at 5Hz. Since Spacemouse-based teleoperation requires considerable expertise, we allocate approximately 5 hours for practice with this system.
Tasks. We focus on insertion tasks that exemplify the challenge of many contact-rich manipulation tasks. We 3D-print a cylinder and an insertion hole base and set up the same task environments for both data collection settings. An example has been shown in Fig. 1, and the dimensions of each piece, including those used in the generalization tasks (Sec. 4.3), are detailed in Appendix. F.
@@ 4.1 Human Tactile Demonstration Collection System
The results in Table 1 support our insights: i) human tactile demonstrations can be collected significantly more efficiently than teleoperation systems for contact-rich tasks, and ii) human tactile demonstrations can seamlessly
integrate human’s tactile feedback and motion capability, whereas teleoperation systems struggle to capture such dynamic tactile-guided motions. These factors together result in much lower data collection efficiency and success rates of the teleoperation systems.
@@ 4.2.1 MimicTouch effectively learns from human tactile demonstrations
In this subsection, we evaluate offline policies learned from human tactile demonstrations using both a nearest-neighbor and a parametric imitation learning method. Our goal is to test if these policies can perform within the desired error tolerance in real-world settings, which is crucial for physical robot execution. To evaluate it, we use our data collection system (Sec. 3.1) to collect 20 noise-free data sequences as the datasets to learn both offline policies. For testing, we gather another 5 data sequences with random noise to emulate the real-world environments (details are in Appendix. G), and compute the mean square error (MSE, defined in Appendix. H) on the testing set. Also, we conduct the real-world experiments and compare the task success rate over 25 policy rollouts.
@@
Additionally, we conduct ablation study on different sensor inputs in Appendix. J. The results suggest that multi-modal tactile feedback is crucial for the success of contact-rich insertion tasks.
@@ 4.2.2 Human tactile demonstrations trains better policies than teleoperated demonstrations
The task success rate for offline rollouts from human tactile demonstration is 40% (10/25), whereas the success rate from Spacemouse-based teleoperation rollouts is only 12% (3/25), and it from Hand-guided teleoperation is 28% (7/25). Then, in the bottom part of Fig. 3, we show the mean and variance of the action serial numbers of three successful rollout trajectories per policy. Human tactile demonstrations show a linear trend with minimal variance, while teleoperation policies display more nonlinearity and higher variance, especially during the insertion phase. This performance discrepancy arises because the majority of contacts occur during this phase, and the teleoperation systems which lack of human tactile feedback is not well-suited for capturing these contact-rich events.
@@
A similar conclusion can be drawn from the screenshots (top part of Fig. 3) for one successful rollout of all policies. Notably, the Human Tactile Demonstrations policy exhibits dynamic tilting for object insertion, which emulates the human’s tactile-guided control strategy (see Fig. 1).
@@ 4.3 Online Policy Improvement and Generalization to New Settings
In this section, we evaluate the final policy trained through online RL. To ensure the robustness of the online policy, we perform domain randomization on the robot’s starting state so that the initial object-hole contact is located differently. For the policy update, at each iteration, we use five newly collected trajectories along with another five randomly selected trajectories from the replay buffer.
@@
MimicTouch policy exhibits superior zero-shot generalization capability. We evaluate the zero-shot generalizability of the MimicTouch final policy. We consider the following generalization settings: i). Shifting Positions (Shift): an insertion task with the hole shifting for 0.8cm in either ±$\pm x$ or ±$\pm y$ directions, ii). Tilting Angles (10°and 20°): an insertion task with the hole angle tilted for 10°or 20°, iii). Two-stage Dense Packing (Two-stage): a two-stage dense insertion task requires the robot to perform two consecutive alignments, iv). Multi-material Dense Packing (Rigid and Soft): an insertion task where the box contains multiple objects, such as a pen, tissues, or a sponge, and v) Furniture Assembly (Assem(I)) and Assem) [53, 54]: an insertion task requires the robot to insert and thread the leg into a small hole of a table. Each setting is depicted in Fig. 5. To demonstrate the complexity of these tasks, we introduce a baseline: Openloop Policy, where we collect five successful insertion trajectories from the initial setting and execute each of these trajectories five times for each of the tasks. See Appendix. L for the details of each task and the policy evaluation process.
@@
Table 2: Task success rate for each generalization task.